# Generating Plausible Distractors for Multiple-Choice Questions via Student Choice Prediction

### v1

@@ 2.1 Distractor Generation
Previous studies on distractor generation can be categorized based on the item format and domain.
Passage-Based
Many studies focus on generating distractors based on passage (e.g., reading material), often using datasets like RACE (Lai et al., 2017), DREAM (Sun et al., 2019), SciQ (Welbl et al., 2017), and Wikipedia (Le Berre et al., 2022). Qiu et al. (2020) proposed the EDGE framework, which reformulates passages and questions through attention mechanisms to generate distractors. Qu et al. (2024) introduced a dual-task training approach where separate training was conducted using passages and questions as input to generate both answers and distractors.
However, as our study focuses on MCQs in the CS domain without relying on passages, these prior works are not directly comparable to ours.
Cloze-Style
Cloze-style formats are commonly used in literacy tests that ask for appropriate words to fill in blanks (Chiang et al., 2022) or in quizzes assessing science knowledge (Ren and Q. Zhu, 2021). Wang et al. (2023) proposed a pseudo Kullback-Leibler Divergence method to regulate distractor generation by considering item discrimination factors. Yu et al. (2024) used a knowledge graph to generate distractors by retrieving relevant triplets and selecting those most aligned with the QA context.
Our framework is not limited to cloze-style questions, which are relatively infrequent in our dataset, and supports a broader range of question types.
@@
Scarlatos et al. (2024) improved the process of generating distractors for math problems by dividing it into two main steps: overgenerate and rank. In the overgenerate phase, they used a large language model (LLM) to generate n distractors, and in the rank phase, a ranker was employed to filter out the top-k distractors most likely to be chosen by students. Feng et al. (2024) explored a kNN-based approach to retrieve in-context examples similar to the target question and used them to generate distractors. Fernandez et al. (2024) proposed the DiVERT, which generates distractors based on learned error representations in math MCQs.
The methods by Scarlatos et al. (2024) and Feng et al. (2024) are used as baselines for comparison with our model. We cannot compare with Fernandez et al. (2024) since their method requires error explanations for each distractor.
@@ 4.3 Distractor Generator
We compare the plausibility of distractors generated by our model, GPT models, the kNN approach proposed by Feng et al. (2024), and human experts (from the base MCQ dataset) as measured by our pairwise ranker (‘DPO, Comb.’ in Table 3). Win/tie/lose counts are calculated per question/distractor in two settings:
@@
We conduct a human evaluation where actual students assess the difficulty of distractors generated by our method. The test comprises 40 MCQs (Python: 20, DB: 10, MLDL: 10). Each question was sampled from the test set of the base MCQ dataset and paired with four distractors, one from each model (SFT, DPO, GPT-3.5-turbo, and GPT-4o), along with a ‘None of the above’ option. The test is taken by 15 college students enrolled in AI courses at our university. [^2] The sample size is larger than the one tested on three individuals in Luo et al. (2024). Based on the selection counts for each distractor, we calculate the plausibility and discrimination index for each model. The discrimination index indicates the ability of each item to differentiate between high- and low-performing students and is calculated as $DI=(U-L)/N$, where $U$ and $L$ denote the number of students in the upper ($U$) and lower ($L$) groups who answered the item correctly, and $N$ is the number of students in each group. More details about the test settings and our rationale are provided in Appendix B.4.
@@ 5.1 Pairwise Ranker
As shown in Table 3, in terms of accuracy, our DPO model achieved an accuracy of 67.5% (row 9), outperforming GPT-3.5-turbo (58.7%, row 1) and GPT-4o (64.0%, row 2) on average. This result is somewhat surprising because our model was trained on reasoning generated by GPT-4o. Moreover, the DPO model significantly outperformed the SFT models (58.7%–65.7%, rows 7–8), particularly in Python, showing the effectiveness of DPO in enhancing the reasoning capability of the model. While Scarlatos et al. (2024)’s method achieved strong performance on math questions in their original work, it exhibited lower accuracy on the CS subjects (48.8%, row 6).
@@
Figure 3 presents survey results from three Master’s students evaluating the reasoning quality of the DPO model on a 5-point Likert scale. The model effectively inferred students’ misconceptions with logical reasoning and accurate knowledge.
@@
We conducted an ablation study to examine the effectiveness of our reasoning method for rank accuracy. As shown in Table 3, for GPT-4o, using our reasoning structure (row 2) substantially outperformed other reasoning methods (rows 3–5), leading us to adopt the current reasoning format for the trained models. Training the model without the reasoning process (row 10) significantly reduced ranking accuracy, highlighting the importance of our reasoning method.
@@ 5.2 Distractor Generator
Table 4 summarizes the win/lose counts of our distractor generators against GPT models, Feng et al. (2024), and human-authored distractors, as evaluated by our pairwise ranker (DPO-based). Our DPO model generated more plausible distractors than baseline models in most cases. Compared to human-authored distractors, our DPO model excelled in Python but underperformed in DB and MLDL. This discrepancy may be due to the underrepresentation of these subjects in our dataset, leading to limited exposure during training.
@@
We additionally evaluated the text similarity between model-generated distractors and human-authored ones, as well as their validity (i.e., whether they are indeed incorrect options for the question). Our model (DPO) showed greater text similarity to human-authored distractors than GPT-3.5-turbo and GPT-4o. It also demonstrated higher validity compared to GPT-3.5-turbo, particularly excelling in questions that ask for incorrect statements. Detailed analyses can be found in Appendix B.5 and B.6.
@@ Appendix A Pairwise Ranker
A.1 Prompt of Pairwise Ranker
The instruction prompts of pariwise ranker are in Table 13 (Reasoning), 14 (Rubric), 15 (G-Eval), 16 (Discussion) and 17 (Scarlatos et al., 2024). We used the same prompt (Reasoning) with GPT models and ours (SFT, DPO).
A.2 Pairwise Ranker SFT and DPO Settings
@@
A.3 GPT Prompt for Making Pairwise Ranker Training Data
The instruction prompt for making pairwise ranker training data is in Table 18. To enhance the diversity of expressions and reasoning used in the samples, two reasoning examples are generated for each pair—one with temperature set to 0 and the other to 1.
A.4 Experiment - Consistency in Rank Prediction
Table 6 demonstrates that our pairwise ranker exhibits relatively robust to positional bias. In comparison to GPT-3.5-turbo, which required an average of more than two attempts to produce consistent results when the input order was altered, our model (DPO) was able to achieve consistent results with significantly fewer attempts. Additionally, our model (DPO) slightly outperformed GPT-4o by requiring fewer average generation attempts.
@@ A.5 Plausibility Factors
We used GPT-4o to summarize and categorize reasoning samples where our pairwise ranker accurately predicted the rankings on the test set, and selected six representative examples per question type. Definitions for each category are in Table 19 (Code Type) and 20 (Statement Type).
@@ A.6 Human Evaluation
Ablation (w/o Reasoning) 2.013 2.034 2.036 2.028
@@ B.1 Prompt for Distractor Generator
The instruction prompt of our distractor generator is in Table 23. We used the same prompt with GPT models and ours. But we instructed to generate in json format for GPT models for stability issue.
@@
B.2 Distractor Generator SFT and DPO Settings
@@ B.3 Experiment - Plausibility
Next, our model exhibited higher plausibility in the statement type compared to GPT-4o. When compared with the validity results in Appendix B.6 (Statement), it can be seen that GPT-4o generated more obvious statements, resulting in a lower risk of invalid distractors but making the difficulty level lower. For the code type, both models generated distractors that were not far from the correct answer. However, in the case of Python, the distractors generated by our model were slightly less plausible than those of GPT-4o, likely because the latter made better use of partial errors in the code.
  [0.1cm] per Question (Win/Lose) [0.1cm] per Distractor (Win/Lose)
@@
Table 9: Plausibility evaluation on the distractor generator, categorized by question type. This table further details the results from Table 4 and 8, Setting B, Ours (DPO).
@@ B.4 Human Evaluation
[…]
@@
Table 10: Text similarity evaluation on distractor generators. sB refers to sBLEU, BS refers to BERTScore.
@@
Table 11: Validity evaluation on distractor generators. Each question can be categorized as either Correct/Incorrect or as Code/Statement. Correct and Incorrect refer to question types such as selecting the correct/incorrect statement, respectively. Code type question indicates cases where the answer (and distractor) is in the form of filling in blanks or matching outputs in code, and Statement refers to cases composed of explanatory statements about a concept. Mistral is a model that has not been fine-tuned, w/o $T$ is the result of using a prompt that generates distractors directly without specifying the distractor type, and w/ $T$ is the result using the same prompt as Ours.
DI
B.5 Additional Evaluation - Text Similarity
Table 10 presents the text similarity evaluation results for the distractor generator. sBLEU [^5] https://github.com/mjpost/sacrebleu and BERTScore [^6] https://github.com/Tiiiger/bert_score were used as the text similarity metrics. For sBLEU, the ‘smooth_method’ was set to ‘exp’, and the default parameters were used for BERTScore. In terms of sBLEU, our model (DPO) generates distractors that are most similar to human-authored ones across the majority of subjects.
B.6 Additional Evaluation - Validity
Validity refers to whether the distractors are indeed incorrect options for the question. Table 11 shows the proportion of valid distractors generated by each model according to the type of question. Our models demonstrate stable validity across various question types (e.g., Correct/Incorrect, Code/Statement), significantly outperforming GPT-3.5-turbo and pre-trained Mistral. This highlights the importance of the proposed methodology—first generating the type ($T$) such as ‘Correct/Incorrect knowledge’—in enhancing validity.
Studies have shown that LLMs perform poorly on tasks involving negation (Varshney et al., 2024), and in a similar vein, GPT-3.5-turbo and Mistral show significantly lower validity when generating distractors for question types that require selecting an incorrect option (in Incorrect type, the distractors should actually represent correct knowledge, but these models mostly generated distractors with incorrect knowledge). However, after going through SFT and DPO, the proportion of valid distractors generated for such types greatly increases, indicating that the proposed methodology in this study (first generating types such as ‘Correct/Incorrect knowledge’) plays an important role in improving validity. Meanwhile, there is a slight decrease in validity after DPO compared to SFT, which appears to be a trade-off arising from the process of creating more confusing distractors.
  [0.1cm] per Question (Win/Tie/Lose) [0.1cm] per Distractor (Win/Lose)
@@ B.6 Additional Evaluation - Validity
B.7 Ablation Study
The training settings used for the ablation study are identical to those of our distractor generator training setup (Appendix B.2), except that the base MCQ dataset was used as the training data instead of the student choice dataset. Table 12 presents the results of the ablation study (compare with Table 4 and 8).
B.8 Error Analysis

### v2

@@ 2.1 Distractor Generation
Previous studies on distractor generation can be categorized according to the question format and domain.
Passage-Based Format
This format is used for exams that evaluate accurate knowledge based on provided textual content, with datasets such as RACE (Lai et al., 2017), DREAM (Sun et al., 2019), SciQ (Welbl et al., 2017), and Wikipedia commonly used to generate MCQs (Le Berre et al., 2022). As a distractor generation model for this format, Offerijns et al. (2020) fine-tuned GPT-2 and ensured the validity of MCQs through an external QA filtering step. Qiu et al. (2020) proposed the EDGE framework, which reformulates passages and questions through attention mechanisms to generate distractors. Qu et al. (2024) introduced a dual-task training approach in which separate training was conducted using passages and questions as input to generate both answers and distractors.
However, since our study focuses on MCQs in the CS domain without relying on passages, these prior works are not directly comparable to ours.
Cloze-Style Format
This format is commonly used in literacy tests and science quizzes, where test-takers fill in blanks with appropriate words (Chiang et al., 2022; Ren and Q. Zhu, 2021). Wang et al. (2023) proposed a pseudo Kullback-Leibler Divergence method to regulate distractor generation by considering item discrimination factors. Yu et al. (2024) used a knowledge graph to generate distractors by retrieving relevant triplets and selecting those most aligned with the QA context.
Our framework is not limited to cloze-style questions, which are relatively rare in our dataset, and instead supports a broader range of question types.
@@
Scarlatos et al. (2024) improved the process of generating distractors for math problems by dividing it into two main steps: overgenerate-and-rank. In the overgeneration phase, they used a large language model (LLM) to generate n distractors, while in the ranking phase, a ranker was employed to select the top-k distractors most likely to be chosen by students. Feng et al. (2024) explored a kNN-based approach to retrieve in-context examples similar to the target question and used them to generate distractors. Fernandez et al. (2024) proposed the DiVERT, which generates distractors based on learned error representations in math MCQs. Hang et al. (2024) utilize retrieval-augmented generation (RAG) and chain-of-thought (CoT) for generating relevant and challenging MCQs.
We use the methods by Scarlatos et al. (2024), Feng et al. (2024) and Hang et al. (2024) as baselines for comparison with our model. We cannot compare with Fernandez et al. (2024) since their method requires error explanations for each distractor.
@@ 4.3 Distractor Generator
We compare the plausibility of distractors generated by our model, GPT models, a kNN approach (Feng et al., 2024), a CoT prompting approach (Hang et al., 2024), and human experts (from the base MCQ dataset) as measured by our pairwise ranker (‘DPO, Comb.’ in Table 3). Win/tie/lose counts are calculated per question/distractor in two settings:
@@
We conduct a human evaluation where actual students assess the difficulty of distractors generated by our method. The test comprises 40 MCQs (Python: 20, DB: 10, MLDL: 10). Each question was sampled from the test set of the base MCQ dataset and paired with four distractors, one from each model (SFT, DPO, GPT-3.5-turbo, and GPT-4o), along with a ‘None of the above’ option. The test is taken by 15 college students enrolled in AI courses at our university [^2] The sample size is larger than the one tested on three individuals in Luo et al. (2024).. Based on the selection counts for each distractor, we calculate the plausibility and discrimination index for each model. The discrimination index indicates the ability of each item to differentiate between high- and low-performing students and is calculated as $DI=(U-L)/N$, where $U$ and $L$ denote the number of students in the upper ($U$) and lower ($L$) groups who answered the item correctly, and $N$ is the number of students in each group.
We also evaluated the clarity (i.e., whether each distractor is clearly written without ambiguity) and answerability (i.e., whether a student with relevant knowledge can reasonably answer the question, as defined by Moon et al. (2022)) of MCQs composed solely of the distractors generated by our DPO model with 11 Master’s students in data science. More details about the human evaluation are provided in Appendix B.4.
@@ 5.1 Pairwise Ranker
As shown in Table 3, in terms of accuracy, our DPO model achieved an accuracy of 67.5% (row 10), outperforming GPT-3.5-turbo (58.7%, row 1) and GPT-4o (64.0%, row 2) on average. This result is somewhat surprising because our model was trained on reasoning generated by GPT-4o. Moreover, the DPO model significantly outperformed the SFT models (58.7%–65.7%, rows 8–9), particularly in Python, showing the effectiveness of DPO in enhancing the reasoning capability of the model. While Scarlatos et al. (2024)’s method achieved strong performance on math questions in their original work, it exhibited lower accuracy on the CS subjects (48.8%, row 7).
@@
Figure 3 presents survey results from three Master’s students evaluating the reasoning quality of the DPO model on a 5-point Likert scale. These results provide mild to moderate evidence supporting the model’s ability to infer students’ misconceptions through logical reasoning and accurate knowledge.
@@
We conducted an ablation study to examine the effectiveness of our reasoning method for rank accuracy. As shown in Table 3, for GPT-4o, using our reasoning structure (row 2) substantially outperformed other reasoning methods (rows 4–6), leading us to adopt the current reasoning format for the trained models. Training the model without the reasoning process (row 11) significantly reduced ranking accuracy, highlighting the importance of our reasoning method.
@@ 5.2 Distractor Generator
Table 4 summarizes the win/lose counts of our distractor generators against GPT models, Feng et al. (2024), Hang et al. (2024) and human-authored distractors, as evaluated by our pairwise ranker (DPO-based). Our DPO model generated more plausible distractors than baseline models in most cases. Compared to human-authored distractors, our DPO model excelled in Python but underperformed in DB and MLDL. This discrepancy may be due to the underrepresentation of these subjects in our dataset, leading to limited exposure during training.
@@
Expert evaluation of our DPO model on clarity and answerability using a 5-point Likert scale showed that all metrics scored above 4, confirming that most distractors were clear enough to answer the question. Detailed results are provided in Appendix B.4.
@@
We additionally evaluated the similarity between model-generated distractors and human-authored ones, as well as their validity. Our DPO model showed greater text similarity to human-authored distractors than GPT-3.5-turbo and GPT-4o. It also demonstrated higher validity compared to GPT-3.5-turbo, particularly excelling in questions that ask for incorrect statements.
Furthermore, we examined the similarity between model-generated distractors and the correct answer to assess the potential issue of distractors being too similar to the correct answer. Our analysis found no evidence that our models pose a particularly high risk to students because of this issue. Detailed analyses can be found in Appendix B.5.
@@
[…]
Setting A Setting B
Ours (SFT) Ours (DPO) Ours (SFT) Ours (DPO)
Python GPT-3.5 71/65 81/56 100/74 105/59
GPT-4o 57/78 74/60 87/53 95/43
DB GPT-3.5 89/67 88/51 95/64 107/54
GPT-4o 93/67 90/53 103/47 120/33
MLDL GPT-3.5 59/93 59/82 71/108 83/96
GPT-4o 57/100 67/80 77/98 90/90
English GPT-3.5 44/44 41/42 78/65 83/61
GPT-4o 44/46 39/46 80/64 84/59
Table 6: Plausibility evaluation of distractor generators on four publicly available datasets (GPT-generated CS questions and a Korean high school English exam). For the English questions, plausibility was evaluated using GPT-4o due to its higher performance. Win/lose counts of our models (columns) against baselines (rows), averaged over two evaluations.
  Rank Accuracy $\uparrow$
GPT-3.5 GPT-4o Ours (SFT)
English 0.483 0.608 0.573
Table 7: Evaluation results on pairwise rankers for English questions. The results were averaged over five generations for each model.
(5) Generalizability
To verify the generalizability of our approach beyond the base MCQ dataset and the CS domain, we conducted additional experiments on two publicly available datasets: (1) newly generated CS questions created using GPT-4o and (2) high school English exam questions.
For CS questions, we generated 100 MCQs per subject using GPT-4o and built a new student choice dataset to train a distractor generator. The results in Table 6, evaluated using our pairwise ranker, are consistent with those from the base MCQ dataset, reaffirming that plausibility improves with DPO over SFT.
For English questions, we used 88 questions from a South Korean high school exam dataset to train a pairwise ranker and a distractor generator. In Table 7, our pairwise ranker, despite limited training data, outperformed GPT-3.5-turbo and closely approached GPT-4o. Similarly, Table 6 shows that in Setting B, where more distractors were compared, our DPO model achieved higher plausibility than GPT models, reflecting the trends observed in CS subjects.
@@ Appendix A Pairwise Ranker
A.1 Prompt for Pairwise Ranker
The instruction prompts for the pariwise ranker are in Table 17 (Reasoning), 18 (Rubric), 19 (G-Eval), 20 (Discussion) and 21 (Scarlatos et al., 2024). We used the same prompt (Reasoning) with GPT models and ours (SFT, DPO).
A.2 SFT and DPO Settings for Pairwise Ranker
@@
A.3 Prompt for Generating Pairwise Ranker Training Data
The instruction prompt for generating pairwise ranker training data is in Table 22. To enhance the diversity of expressions and reasoning used in the samples, two reasoning examples are generated for each pair—one with temperature set to 0 and the other to 1—using GPT-4o.
A.4 Consistency in Rank Prediction
Table 8 demonstrates that our pairwise ranker exhibits relatively robust to positional bias. In comparison to GPT-3.5-turbo, which required an average of more than two attempts to produce consistent results when the input order was altered, our DPO model was able to achieve consistent results with significantly fewer attempts. Additionally, our DPO model slightly outperformed GPT-4o by requiring fewer average generation attempts.
@@ A.5 Plausibility Factors
We used GPT-4o to summarize and categorize reasoning samples where our pairwise ranker accurately predicted the rankings on the test set, and selected six representative examples per question type. Definitions for each category are in Table 23 (Code Type) and 24 (Statement Type).
@@ A.6 Human Evaluation
Ours (SFT w/o Reasoning) 2.013 2.034 2.036 2.028
Table 8: Consistency evaluation results on pairwise rankers. The results were averaged over five generations for each model.
@@ B.1 Prompt for Distractor Generator
The instruction prompt for our distractor generator is in Table 27. We used the same prompt for both GPT models and ours. However, we instructed the GPT models to generate outputs in JSON format for stability reasons.
@@
B.2 SFT and DPO Settings for Distractor Generator
@@ B.3 Plausibility Evaluation
Next, our model exhibited higher plausibility in the statement type compared to GPT-4o. When compared with the validity results in Appendix B.5 (‘Statement’), it can be seen that GPT-4o generated more obvious statements, resulting in a lower risk of invalid distractors but making the difficulty level lower. For the code type, both models generated distractors that were not far from the correct answer. However, in the case of Python, the distractors generated by our model were slightly less plausible than those of GPT-4o, likely because the latter made better use of partial errors in the code.
[…]
@@
Table 11: Plausibility evaluation on the distractor generator, categorized by question type. This table further details the results from Table 4 and 10, Setting B, Ours (DPO).
@@ B.4 Human Evaluation
  5-point Likert Scale $\uparrow$
Python DB MLDL
Clairity 4.418 4.282 4.373
Answerability 4.414 4.264 4.382
Table 12: Evaluation results on the clarity and answerability of distractors generated by our DPO model.
  sBLEU $\uparrow$ BERTScore $\uparrow$
@@
Table 13: Similarity between model-generated and human-authored distractors.
  Similarity
Human GPT-3.5 GPT-4o Feng et al. (2024) Hang et al. (2024) Ours (SFT) Ours (DPO)
Python 0.53 (0.04) 0.54 (0.05) 0.60 (0.05) 0.58 (0.04) 0.59 (0.05) 0.57 (0.06) 0.59 (0.06)
DB 0.55 (0.05) 0.55 (0.05) 0.60 (0.05) 0.58 (0.04) 0.59 (0.05) 0.59 (0.06) 0.61 (0.06)
MLDL 0.53 (0.04) 0.47 (0.05) 0.51 (0.05) 0.52 (0.04) 0.52 (0.05) 0.50 (0.05) 0.53 (0.05)
Table 14: Similarity between model-generated distractors and correct answers. Numbers in parentheses represent variance.
DI
Clairity and Answerability
We concern that excessively ambiguous distractors could hinder educational assessment. To verify whether such issues exist, we asked 11 experts to evaluate the clarity and answerability of distractors generated by our DPO model on a 5-point Likert scale. The results, averaged across a total of 40 questions (Python: 20, DB: 10, MLDL: 10) for each subject, are in Table 12. These results confirm that the potential side effects of high plausibility, which were a concern, do not appear to be present in our model.
B.5 Additional Evaluations
Similarity Between Model-Generated and Human-Authored Distractors
Table 13 presents the similarity evaluation results for the distractor generator. sBLEU [^5] https://github.com/mjpost/sacrebleu and BERTScore [^6] https://github.com/Tiiiger/bert_score were used as the text similarity metrics. For sBLEU, the ‘smooth_method’ was set to ‘exp’, and the default parameters were used for BERTScore. In terms of sBLEU, our model (DPO) generates distractors that are most similar to human-authored ones across the majority of subjects.
Similarity Between Model-Generated Distractors and Correct Answers
We evaluated the semantic similarity between distractors and correct answers using OpenAI’s text-embedding-3-small model. Cosine similarity scores were computed between correct answers and distractors generated by each model.
As shown in Table 14, similarity scores are largely consistent across models, with no evidence that our SFT or DPO models generate distractors that are excessively similar to correct answers. For instance, the average similarity score of human-authored distractors (0.53) is comparable to that of our DPO model (0.59) for Python.
  • Question: "Choose the incorrect statement."
  • Correct Answer: "If an if statement’s condition is False, its content executes before the else statement."
  • Distractor 1 (similarity = 0.59): "An if-else statement can include multiple conditions using elif in some languages."
  • Distractor 2 (similarity = 0.53): "An if statement can exist without an else statement in most languages."
A 0.06 difference in similarity score does not indicate a meaningful impact on quality, as the first distractor is not noticeably more similar to the correct answer than the second.
  Validity $\uparrow$
@@ B.5 Additional Evaluations
Table 15: Validity evaluation on distractor generators. Mistral is a model that has not been fine-tuned, w/o $T$ is the result of using a prompt that generates distractors directly without specifying the distractor type, and w/ $T$ is the result using the same prompt as Ours.
Validity
Validity refers to whether the distractors are indeed incorrect options for the question. We assessed validity by categorizing questions as either Correct/Incorrect or Code/Statement. ‘Correct’ and ‘Incorrect’ refer to question types where the task is to select the correct or incorrect statement, respectively. ‘Code’ type questions involve cases where the answer (and distractors) take the form of filling in blanks or matching outputs in code. ‘Statement’ refers to questions composed of explanatory statements about a concept.
Table 15 shows the proportion of valid distractors generated by each model according to the type of question. Our models demonstrate stable validity across various question types, significantly outperforming GPT-3.5-turbo and pre-trained Mistral. This highlights the importance of the proposed methodology—first generating the type ($T$) such as ‘Correct/Incorrect knowledge’—in enhancing validity.
Studies have shown that LLMs perform poorly on tasks involving negation (Varshney et al., 2024), and in a similar vein, GPT-3.5-turbo and Mistral show significantly lower validity when generating distractors for question types that require selecting an incorrect option (in ‘Incorrect’ type, the distractors should actually represent correct knowledge, but these models mostly generated distractors with incorrect knowledge). However, after going through SFT and DPO, the proportion of valid distractors generated for such types greatly increases, indicating that the proposed methodology in this study (first generating types such as ‘Correct/Incorrect knowledge’) plays an important role in improving validity. Meanwhile, there is a slight decrease in validity after DPO compared to SFT, which appears to be a trade-off arising from the process of creating more confusing distractors.
[…]
@@
B.6 Ablation Study
The training settings used for the ablation study are identical to those of our distractor generator training setup (Appendix B.2), except that the base MCQ dataset was used as the training data instead of the student choice dataset. Table 16 presents the results of the ablation study (compare with Table 4 and 10).
B.7 Error Analysis